OpenStudy (anonymous):

Calculate the energy in J of 0.32 moles of photons whose frequency is 2.6 x 10^15 ?

OpenStudy (anonymous):

You must know the formula for calculating the energy... in this situation. Do you know the formula?

OpenStudy (aaronq):

E=hv h is planks constant v= frequency this gives you the energy of 1 photon so multiply this by the amount of photons you have: amount of photons = avogadros constant x number of moles

OpenStudy (anonymous):

that v is actually "new (pronounced as new)"

OpenStudy (anonymous):

So @jennilalala with the help of : E = hv formula, can you find the energy by putting the values given ?

OpenStudy (anonymous):

"v" --> \(\large{\color{blue}{\nu}}\) .... right @aaronq ?

OpenStudy (anonymous):

I am back sorry ahaah. yes i know the formulas. i just dont know what to do if it says ".32 moles of photon.."

OpenStudy (anonymous):

Do you the planck's constant ?

OpenStudy (anonymous):

6.626x10^23 ?? not by heart...yet...

OpenStudy (aaronq):

haha yep, v= "new"

OpenStudy (anonymous):

10^-34

OpenStudy (anonymous):

OK so it is \[\large { h = 6.626 \times 10^{-34} m^{-1}}\]

OpenStudy (anonymous):

oh wait it is m^2 kg/s

OpenStudy (anonymous):

So right, now solve this : \[\large{E = 6.626 \times 10^{-34} \times 2.6 \times 10^{15}}\] What do you get?

OpenStudy (anonymous):

1.73x10^-18

OpenStudy (anonymous):

My calculator says it as : \(\large{\color{blue}{ 17.2276 * 10^{-19}}}\) that is 1.72276 * 10^{-18}

OpenStudy (anonymous):

OK so now .. this is the energy for \(1\) photon ... now we to calculate for 0.32 moles of photons.

OpenStudy (anonymous):

See ... no. of photons = Avogadros's constant * no. of moles of photons...

OpenStudy (anonymous):

You know avogadro's constant?

OpenStudy (anonymous):

6.022x10^23

OpenStudy (anonymous):

Good! So now, solve this one : \[\large{\color{red}{\mathbb{No. of Photons}} = \color{blue}{6.022 * {10}^{23}} \times \color{green}{0.32}}\]

OpenStudy (anonymous):

1.93x10^23

OpenStudy (anonymous):

then i multiply the two calculations?

OpenStudy (anonymous):

Correct! So this is your No. of photons. Now multiply this result with "1.73 * 10^{-18} "

OpenStudy (anonymous):

3.3x10^-5

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

not negative typo lol

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

So finally you have your correct way, I hope you got it ?

OpenStudy (anonymous):

Yes thank you so much!

OpenStudy (anonymous):

You're welcome @jennilalala . keep up the good work.